The concentration of a Fe2+ solution is determined by titrating it with a 0.1775 M solution of permanganate. The balanced net ionic equation for the reaction is shown below.

MnO4-(aq) + 5 Fe2+(aq)+8 H3O+(aq) Mn2+(aq) + 5 Fe3+(aq)+12 H2O(l)

In one experiment, 16.84 mL of the 0.1775 M MnO4- solution is required to react completely with 25.00 mL of the Fe2+ solution. Calculate the concentration of the Fe2+ solution.

______ M
1
Expert's answer
2019-03-12T08:55:14-0400

Solution.

MnO4-(aq) + 5 Fe2+(aq)+8 H3O+(aq) Mn2+(aq) + 5 Fe3+(aq)+12 H2O(l)

Firstly, we write the equation of the law of equivalents during titration:

C(KMnO4)*V(KMnO4) = C(Fe2+)*V(Fe2+)

C(Fe2+) = (C(KMnO4)*V(KMnO4))/V(Fe2+)

C(Fe2+) = (0.1775 * 16.84)/25.00 = 0.1196 M

Answer:

C(Fe2+) = (0.1775 * 16.84)/25.00 = 0.1196 M